OpenStudy (anonymous):

Consider the sequence: 2,6,18,54 Let n= the term number in the sequence. Let A(n)= the value of the nth term, of the sequence. 1)Complete each statement: a. A(1) = 2 = 2 x 3 b. A(2) = 6 = 2 x 3= 2 x 3 c. A(3) = 18= 2 x 3 x 3= 2 x 3 d. A(4) = 54= 2 x 3 x 3 x 3= 2 x 3 What is the relationship between the exponent of the base 3 and the value of n?

OpenStudy (anonymous):

It is a geometric series ratio is 3

OpenStudy (anonymous):

i have given enough hints

OpenStudy (anonymous):

hope you can work out

OpenStudy (campbell_st):

the exponent is (n - 1) e.g. 2 = 2 x 3^(1 -1) = 2 x 3^0 = 2

OpenStudy (anonymous):

Thanks For The Help But I Alrediee Knew The Ratio.

OpenStudy (anonymous):

O.O That Makes Sense campbell.!

OpenStudy (anonymous):

Thank Youu Veriee Much.! :D